# Geometric Bounds for Persistence

### v1

@@ 3.4 Robust Bounds for Čech Lifespans via Treewidth
The estimates above have a drawback: There is no a priori way to relate the width of $\mathcal{X}\subset\mathbf{Y}$ and the width of its $r$-neighborhood $\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})$, while the bounds in Corollaries 3.121 and 1 depend on the width of a certain neighborhood of $\mathcal{X}$. By narrowing down the class of potential cores even further we can guarantee some bounds. For example, this happens if the core is an affine $k$-dimensional subspace (see Section 3.5). Here we introduce a modified version of treewidth that allows for bounds on lifespans in terms of $\mathcal{X}$ itself and not of its neighborhood. We start with an auxiliary definition that we will use to restrict the geometric complexity of potential cores.
@@
Recall that a retraction $f:\mathbf{Y}\to\mathcal{T}$ is a continuous map fixing every point of $\mathcal{T}$. Since $\mathbf{Y}$ is contractible, the existence of a retraction implies that $\mathcal{T}$ is contractible.
@@
    If $\mathcal{T}$ is nonempty closed convex subset of $\mathbf{Y}$, and $f$ is a continuous nearest-point projection (metric projection) on $\mathcal{T}$, then $f$ is 00-robust. [^10] It is known (see the discussion after [Meg12, Corollary 5.1.19]) that a nearest-point projection exists if $\mathbf{Y}$ is reflexive; but it does not have to be continuous in general.
@@
For any compact set $\mathcal{X}$ sitting in a Banach space $\mathbf{Y}$, any integer $k\geq 0$, any real numbers $C,r\geq 0$, and any homology class $\omega\in\operatorname{\tilde{H}}_{k}(\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y}))$,
@@
Let our homology class $\omega\in\operatorname{\tilde{H}}_{k}(\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y}))$ be represented by a singular cycle $S_{0}$ supported in $\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})$. Since $S_{0}$ is compact as a subset of $\mathbf{Y}$, it can be covered by finitely many open balls $B_{r}(x),x\in\mathcal{X}^{\prime}\subset\mathcal{X},|\mathcal{X}^{\prime}|<\infty$ (here $B_{r}(x)$ denotes the open ball of radius $r$ centered at $x$). Consider the union $U=\bigcup\limits_{x\in\mathcal{X}^{\prime}}B_{r}(x)$ and choose any partition of unity subordinate to the cover of $U$ by $B_{r}(x),x\in\mathcal{X}^{\prime}$. This partition of unity induces a map from $U$ to the nerve $Z$ of that cover. By the nerve theorem this map $\phi:U\to Z$ is a homotopy equivalence, with an evident homotopy inverse $\psi:Z\to U$, which sends every vertex of $Z$ to the center of the corresponding ball in the cover $U=\bigcup\limits_{x\in\mathcal{X}^{\prime}}B_{r}(x)$, and extends affinely on the rest of $Z$. Therefore, $\omega|_{U}=\psi_{*}\phi_{*}\omega|_{U}\in\operatorname{\tilde{H}}_{k}(U)$, and the class $\omega$ can be represented by a cycle $S$ supported in $\psi(Z)$; in other words, $S$ is obtained by gluing flat $k$-simplices with vertices in $\mathcal{X}$, and with the circumradius of each simplex at most $r$.
@@
By the robustness property of $f$, every point $x\in S$ is moved by a distance at most $\delta+Cr$, so the continuous deformation always stays in $\mathcal{N}_{\delta+Cr}(S\subset\mathbf{Y})$. At time $t=1$, the deformed cycle $h(S,1)$ lies in $\mathcal{T}$ and bounds a $(k+1)$ chain within $h(S,1)$, because $\mathcal{T}$ is of dimension $k$ with trivial $\operatorname{\tilde{H}}_{k}(\mathcal{T})$. Therefore, $S$ is nullhomologous in $\mathcal{N}_{\delta+Cr}(S\subset\mathbf{Y})$ (basically, $S$ bounds the $(k+1)$-dimensional trace of the homotopy $h$). ∎
@@
We again employ formula (2.3), now bounding filling radii by 3.20:
@@
For $b_{\omega}\geq 1$, if we divide the entire inequality by $b_{\omega}$, we get the result. ∎
Remark 3.21 (Multiplicative Persistence).
The corollary above implies a robust estimate on the multiplicative persistence (see [BKS17, AC22]) of intervals in the persistence barcode of the Čech filtration $\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y})$. Notice that on the right-hand side, there is no reference to the birth time $b_{\omega}$. Therefore, $C$-robust treewidth gives a strong bound for late-born ($b_{\omega}\geq 1$) homology classes in terms of multiplicative persistence. Note also that the proof above also gives the following estimates for $(b_{\omega},d_{\omega})$:
@@
3.5 Robust Bounds for Čech Lifespans via Kolmogorov Width and ${}_{\infty}$
For a given compact set $\mathcal{X}$ in a Banach space $\mathbf{Y}$ (e.g., a point cloud in $\mathbb{R}^{N}$), the usual Principal Component Analysis (PCA) approximates $\mathcal{X}$ by affine $k$-planes through minimizing the $\ell^{2}$-distance from $\mathcal{X}$ to a plane. Here we consider an $\ell^{\infty}$-version of the classical procedure, by optimizing the $\ell^{\infty}$-distance instead. Recall that in the original PCA, the eigenvalues of the covariance matrix are called variances. This motivates the following definition.
Definition 3.22 (${}_{\infty}$).
@@ 3.5 Robust Bounds for Čech Lifespans via Kolmogorov Width and ${}_{\infty}$
$\nu_{k+1}(\mathcal{X}\subset\mathbf{Y})=\inf_{\mathcal{P}\in\mathcal{A}_{k}}\sup_{x\in\mathcal{X}}\mathbf{d}_{\mathbf{Y}}(x,\mathcal{P})$
where $\mathbf{d}_{\mathbf{Y}}(x,\mathcal{P})$ is the distance from $x$ to the $k$-subspace $\mathcal{P}$, i.e., $\mathbf{d}_{\mathbf{Y}}(x,\mathcal{P})=\inf_{p\in\mathcal{P}}\|x-p\|$.
Remark 3.23 (Kolmogorov Width).
@@
The $\infty$-variance is just a different name for Kolmogorov widths (cf. 3.6):
$\nu_{k+1}(\mathcal{X}\subset Y)=\operatorname{KW}_{k}(\mathcal{X}\subset\mathbf{Y}).$
@@
Let $\mathcal{X}$ be a subset of a Banach space $\mathbf{Y}$. Then for any homology class $\omega\in\operatorname{\tilde{H}}_{k}(\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y}))$, $k\geq 0$,
@@
Notice that the corollary above gives strong estimates for point clouds in $\mathbb{R}^{N}$. For example, if the variance $\nu_{k_{0}+1}(\mathcal{X})$ is small, then the corollary above implies that one can ignore homology classes in $\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$ for $k\geq k_{0}$. We should note that, by its very definition, ${}_{\infty}$ is not as robust as the original PCA against outliers. We discuss this difference and pose a statistical question about the relationship between original PCA and persistent homology in Section 6.
@@
Notice that all lifespan bounds in Corollaries 3.12, 1, 1, 1, and 1 are stated in terms of homological lifespans ($d_{\omega}-b_{\omega}$). 2.22 implies that all these bounds therefore apply to the length of every interval in either $\mathrm{PD}_{k}(\mathcal{X})$ or $\widecheck{\mathrm{PD}}_{k}(\mathcal{X}\subset\mathbf{Y}).$
@@ 4 Bounding Lifespans by Spread
This section is concerned with global upper bounds for Čech lifespans of homology classes in arbitrary degrees dimension. Our results in this section are somewhat parallel to the following result for VR lifespans.
Theorem 4.1 (VR-lifespans via Spread; [LMO23, Proposition 9.3 and Remark 9.8]).
@@
Here Katz’s notion of spread is used [Kat83]. By definition, $\operatorname{spread}(\mathcal{X})$ is the infimum of the real numbers $\delta\geq 0$ for which there is a finite subset $A=\{a_{1},\ldots,a_{m}\}\subset\mathcal{X}$ with $\operatorname{diam}(A)\leq\delta$ and $\mathbf{d}^{\mathcal{X}}_{\mathrm{H}}(\mathcal{X},A)\leq\delta$.
In the rest of this section, we work in a Banach space $(\mathbf{Y},\|\cdot\|)$ and, for each compact set $\mathcal{X}\subset\mathbf{Y}$, we discuss a certain variation of the spread for which we prove an analogue of 4.1.
@@
A set $\mathcal{T}$, sitting in a Banach space $(\mathbf{Y},\|\cdot\|)$, will be called $\delta$-übercontractible if its neighborhood $\mathcal{N}_{r}(\mathcal{T}\subset\mathbf{Y})$ is contractible for every $r\geq\delta$.
@@
For our purposes we actually need a slightly weaker condition: all neighborhoods $\mathcal{N}_{r}(\mathcal{T})$, $r\geq\delta$, are acyclic (have trivial homology groups) rather than contractible (have trivial homotopy groups). The acyclicity condition is equivalent to having almost trivial Čech persistence diagram (trivial beyond death time $\delta$), and is implied by übercontractibility as defined above.
For constructive examples of übercontractible sets, we refer the reader to Appendix B, where we show that the cut-locus of a convex set is typically 00-übercontractible.
@@
For a compact set $\mathcal{X}\subset\mathbf{Y}$, we define its überspread as
@@
Figure 3: The bound $d-b\leq 2\,\ddot{u}$ in Section 1 is sharp; here $\mathbf{Y}=\mathbb{R}^{2}$ and $\ddot{u}=\operatorname{{\ddot{u}}-spread}(\mathcal{X}\subset\mathbb{R}^{2})$.
@@
[…]
@@
$\operatorname{\tilde{H}}_{k}(i_{r,r+2\delta})=\operatorname{\tilde{H}}_{k}(p_{r})\circ\operatorname{\tilde{H}}_{k}(q_{r+\delta})=0$
@@
Let us compare Section 1 with [LMO23, Proposition 9.3]. Suppose we are given a compact metric space $\mathcal{X}$ and a finite subset $A=\{a_{1},\ldots,a_{m}\}\subset\mathcal{X}$ with $\operatorname{diam}(A)\leq\delta$ and $\mathbf{d}^{\mathcal{X}}_{\mathrm{H}}(\mathcal{X},A)\leq\delta$. In other words, the Katz spread of $\mathcal{X}$ does not exceed $\delta$:
@@
Embed $\mathcal{X}$ in $L^{\infty}(\mathcal{X})$ via the Kuratowski map $\kappa$, and note that the set $\kappa(A)\subset L^{\infty}(\mathcal{X})$ is $\delta/2$-übercontractible. Hence, we can apply Section 1 and conclude that Čech lifespans of $\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X})$ do not exceed $2\delta$. Therefore, VR lifespans of $\mathcal{X}$ do not exceed $4\delta$. This is weaker than the conclusion of [LMO23, Proposition 9.3], which tells us that in this situation VR lifespans of $\mathcal{X}$ do not exceed $\operatorname{spread}(\mathcal{X})\leq\delta$.
In general, the estimate in Section 1 cannot be improved (see Figure 3). However, the factor of $2$ in that estimate can sometimes be removed if the comparison set $\mathcal{T}$ admits a $1$-Lipschitz nearest-point projection from $\mathbf{Y}$.
@@
Let $\mathcal{X}$ be a compact subset of a Banach space $(\mathbf{Y},\|\cdot\|)$. Assume that $\mathrm{conv}(\mathcal{X})$ admits a $1$-Lipschitz nearest-point projection to a closed convex set $\mathcal{T}\subset\mathbf{Y}$ [^11] For example, this is true whenever $\mathbf{Y}$ is Euclidean. Then any $\omega\in\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$, $k\geq 0$, has lifespan at most $\mathbf{d}_{\mathrm{H}}^{\mathbf{Y}}(\mathcal{X},\mathcal{T})$.
Proof.
Denote $\delta=\mathbf{d}_{\mathrm{H}}^{\mathbf{Y}}(\mathcal{X},\mathcal{T})$. By assumption, there exists a $1$-Lipschitz nearest-point projection $\pi:\mathrm{conv}(\mathcal{X})\rightarrow\mathcal{T}$; there is also a (possibly discontinuous) map $j:\mathcal{T}\to\mathcal{X}$ such that
@@
[…]
@@
Now, going back to the proof of the theorem: apply the homology functor $\operatorname{\tilde{H}}_{k}$ to the diagram above and obtain the following commutative diagram:
[…]
Now, $\operatorname{\tilde{H}}_{k}(\mathcal{C}_{\epsilon}(\mathcal{T}\subset\mathbf{Y}))=0$ since $\mathcal{T}$ is convex. Then,
$\operatorname{\tilde{H}}_{k}(i_{\epsilon,\epsilon+\delta})=\operatorname{\tilde{H}}_{k}(j_{\epsilon})\circ\operatorname{\tilde{H}}_{k}(\pi_{\epsilon})=0,$

### v2

@@ 3.4 Robust Bounds for Čech Lifespans via Treewidth
The estimates provided in the preceding sections have a drawback: there is no a priori method to directly relate the width of $\mathcal{X}\subset\mathbf{Y}$ to the width of its $r$-neighborhood $\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})$. This poses a challenge, as the bounds in Corollaries 3.12, 1.2, and 1.2 rely on the width of a specific neighborhood of $\mathcal{X}$. However, by further refining the class of potential “cores”, we can ensure the derivation of more effective bounds. For example, this is possible if the core is an affine $k$-dimensional subspace (see Section 3.5).
Here, we introduce a modified version of treewidth that enables bounding lifespans directly in terms of $\mathcal{X}$ itself, rather than its neighborhood. To this end, we introduce an auxiliary definition to help constrain the geometric complexity of potential “cores”.
@@
Recall that, given $\mathcal{T}\subset\mathbf{Y}$, a retraction $f:\mathbf{Y}\to\mathcal{T}$ is a continuous map fixing every point of $\mathcal{T}$. Since $\mathbf{Y}$ is contractible, the existence of a retraction implies that $\mathcal{T}$ is contractible.
@@
    If $\mathcal{T}$ is nonempty closed convex subset of $\mathbf{Y}$, and $f$ is a continuous nearest-point projection (metric projection) on $\mathcal{T}$, then $f$ is 00-robust. [^14] It is known that a nearest-point projection exists, if $\mathbf{Y}$ is reflexive, but it does not have to be continuous in general; see for example the discussion after [Meg12, Corollary 5.1.19].
@@
For any compact set $\mathcal{X}$ sitting in a Banach space $\mathbf{Y}$, any integer $k\geq 0$, any real numbers $C\geq 0$ and $r>0$, and any homology class $\omega\in\widetilde{\mathrm{H}}_{k}(\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y}))$,
@@
Let our homology class $\omega\in\widetilde{\mathrm{H}}_{k}(\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y}))$ be represented by a singular cycle $S_{0}$ supported in $\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})$. Since $S_{0}$ is compact as a subset of $\mathbf{Y}$, it can be covered by finitely many open balls $B_{r}(x),x\in\mathcal{X}^{\prime}\subset\mathcal{X},|\mathcal{X}^{\prime}|<\infty$ (here $B_{r}(x)$ denotes the open ball of radius $r$ centered at $x$). Consider the union $U=\bigcup\limits_{x\in\mathcal{X}^{\prime}}B_{r}(x)$ and choose any partition of unity subordinate to the cover of $U$ by $B_{r}(x),x\in\mathcal{X}^{\prime}$. This partition of unity induces a map from $U$ to the nerve $L$ of that cover. By the nerve theorem this map $\phi:U\to L$ is a homotopy equivalence, with an evident homotopy inverse $\psi:L\to U$, which sends every vertex of $L$ to the center of the corresponding ball in the cover $U=\bigcup\limits_{x\in\mathcal{X}^{\prime}}B_{r}(x)$, and extends affinely on the rest of $L$. Therefore, $\omega|_{U}=\psi_{*}\circ\phi_{*}(\omega|_{U})\in\widetilde{\mathrm{H}}_{k}(U)$, and the class $\omega$ can be represented by a cycle $S$ supported in $\psi(L)$; in other words, $S$ is obtained by gluing flat $k$-simplices with vertices in $\mathcal{X}$, and with the circumradius of each simplex at most $r$.
@@
By the robustness property of $f$, every point $x\in S$ is moved by a distance at most $\delta+Cr$, so the continuous deformation always stays in $\mathcal{N}_{\delta+Cr}(S\subset\mathbf{Y})$. At time $t=1$, the deformed cycle $h(S,1)$ lies in $\mathcal{T}$ and bounds a $(k+1)$ chain within $h(S,1)$, because $\mathcal{T}$ is of dimension $k$ with trivial $\widetilde{\mathrm{H}}_{k}(\mathcal{T})$. Therefore, $S$ is nullhomologous in $\mathcal{N}_{\delta+Cr}(S\subset\mathbf{Y})$ (basically, $S$ bounds the $(k+1)$-dimensional trace of the homotopy $h$). ∎
@@
We again employ formula (2.3), now bounding filling radii by 3.21:
@@
For $b_{\omega}\geq 1$, if we divide the entire inequality by $b_{\omega}$, we obtain the result. ∎
Remark 3.22 (Comments about Section 1.2).
We make the following remarks.
  • Notice that on the right-hand side, there is no reference to the birth time $b_{\omega}$. Therefore, $C$-robust treewidth yields strong control on late-born homology classes (i.e. those satisfying $b_{\omega}\geq 1$).
  • If the the assumption $b_{\omega}\geq 1$ is changed to $b_{\omega}\geq\tfrac{1}{\alpha}$ for some $\alpha>0$ then the conclusion becomes $\tfrac{d_{\omega}}{b_{\omega}}\leq 1+C+\alpha\operatorname{TW}^{C}_{k}(\mathcal{X}\subset\mathbf{Y})$.
  • The corollary implies a robust estimate on the multiplicative persistence of intervals in the persistence barcode of the Čech filtration $\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y})$. See Bobrowski, Kahle and Skraba [BKS17] and Adams and Coskunuzer [AC22] for studies of multiplicative persistence.
  • Additionally, the proof above gives the following estimates for $(b_{\omega},d_{\omega})$:
@@
3.5 Robust Bounds for Čech Lifespans via Kolmogorov Width and PCA
@@
Since its introduction by Pearson in 1901 [Pea01], PCA has become one of the most widely used data analysis techniques, owing to its interpretability and the availability of highly efficient computational algorithms. In this work, we investigate a specific $\ell^{\infty}$-variant of classical PCA, designed for compact subsets of Banach spaces, and examine its connections to persistent homology by analyzing how it constrains the lifespans of homology classes.
Definition 3.23 (PCA).
@@
$\nu_{k+1}(\mathcal{X}\subset\mathbf{Y}):=\inf_{\mathcal{P}\in\mathcal{A}_{k}}\sup_{x\in\mathcal{X}}\mathbf{d}_{\mathbf{Y}}(x,\mathcal{P})$
where $\mathbf{d}_{\mathbf{Y}}(x,\mathcal{P})$ is the distance from $x$ to the $k$-subspace $\mathcal{P}$, i.e., $\mathbf{d}_{\mathbf{Y}}(x,\mathcal{P}):=\inf_{p\in\mathcal{P}}\|x-p\|$.
@@
The term $\ell^{\infty}$-variance is just a different name for Kolmogorov widths (cf. 3.6):
$\nu_{k+1}(\mathcal{X}\subset\mathbf{Y})=\operatorname{KW}_{k}(\mathcal{X}\subset\mathbf{Y}).$
@@
Let $\mathcal{X}$ be a subset of a Banach space $\mathbf{Y}$. Then, for any $r>0$ and any homology class $\omega\in\widetilde{\mathrm{H}}_{k}(\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y}))$, $k\geq 0$,
@@
Notice that the corollary above gives strong estimates for point clouds in $\mathbb{R}^{N}$. For example, if the variance $\nu_{k_{0}+1}(\mathcal{X})$ is small, then, in applications where only long-lived homological features are relevant, the corollary implies that one can ignore homology classes in $\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$ for $k\geq k_{0}$. We should note that, by its very definition, PCA is more sensitive than the original PCA against outliers. We discuss this difference and pose a statistical question about the relationship between the original PCA and persistent homology in Section 6.
@@
Notice that all lifespan bounds in Corollaries 3.12, 1.2, 1.2, 1.2, and 1.2 are stated in terms of homological lifespans ($d_{\omega}-b_{\omega}$). 2.23 implies that all these bounds therefore apply to the length of every interval in either $\mathrm{PD}_{k}(\mathcal{X})$ or $\widecheck{\mathrm{PD}}_{k}(\mathcal{X}\subset\mathbf{Y}).$
@@ 4 Bounding Lifespans by Spread
This section is concerned with global upper bounds for Čech lifespans of homology classes in arbitrary degrees dimension. In terms of the conceptual description laid out on page 1.2, to formulate these results we will choose cores to be übercontractible spaces, as defined below.
Our results in this section are somewhat analogous to the following result concerning VR lifespans.
Theorem 4.1 (VR-lifespans via Spread; [LMO24, Proposition 9.19 and Remark 9.18]).
@@
Here Katz’s notion of spread is used [Kat83]; see also [Wil92, Definition 4]. By definition, $\operatorname{spread}(\mathcal{X})$ is the infimum of the real numbers $\delta\geq 0$ for which there is a finite subset $A=\{a_{1},\ldots,a_{m}\}\subset\mathcal{X}$ with $\operatorname{diam}(A)\leq\delta$ and $\mathbf{d}^{\mathcal{X}}_{\mathrm{H}}(\mathcal{X},A)\leq\delta$.
In the rest of this section, we work in a Banach space $\mathbf{Y}$ and, for each compact set $\mathcal{X}\subset\mathbf{Y}$, we consider a certain variant of the notion spread for which we prove an analogue of 4.1.
@@
A set $\mathcal{T}$, sitting in a Banach space $\mathbf{Y}$, will be said to be $\delta$-übercontractible if its neighborhood $\mathcal{N}_{r}(\mathcal{T}\subset\mathbf{Y})$ is contractible for every $r\geq\delta$.
@@
For our purposes a slightly weaker condition would suffice: all neighborhoods $\mathcal{N}_{r}(\mathcal{T}\subset\mathbf{Y})$, $r\geq\delta$, are acyclic (have trivial homology groups) rather than contractible (have trivial homotopy groups). The acyclicity condition is equivalent to having almost trivial Čech persistence diagram (trivial beyond death time $\delta$), and is implied by übercontractibility as defined above.
@@
For a compact subset $\mathcal{X}$ of a Banach space $\mathbf{Y}$, we define its überspread as
@@
Note that for any point $p\in\mathbf{Y}$ the singleton set $\{p\}$ is $\delta$-übercontractible for all $\delta\geq 0$. In particular, this implies that
$\operatorname{{\ddot{u}}-spread}(\mathcal{X}\subset\mathbf{Y})\leq\mathrm{rad}(\mathcal{X}\subset\mathbf{Y}).$
For additional constructive examples of übercontractible sets, we refer the reader to Appendix B, where we show that the cut-locus of a convex set is typically 00-übercontractible.
@@
Figure 3: The bound $d-b\leq 2\,\ddot{u}$ in Section 1.2 is sharp; here $\mathbf{Y}=\mathbb{R}^{2}$ and $\ddot{u}=\operatorname{{\ddot{u}}-spread}(\mathcal{X}\subset\mathbb{R}^{2})$.
@@
${\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})}$${\mathcal{N}_{r+2\delta}(\mathcal{X}\subset\mathbf{Y})}$${\mathcal{N}_{r+\delta}(\mathcal{T}\subset\mathbf{Y})}$$\scriptstyle{p_{r}}$$\scriptstyle{i_{r,r+2\delta}}$$\scriptstyle{q_{r+\delta}}$
@@
$\widetilde{\mathrm{H}}_{k}(i_{r,r+2\delta})=\widetilde{\mathrm{H}}_{k}(p_{r})\circ\widetilde{\mathrm{H}}_{k}(q_{r+\delta})=0$
@@
Let us compare Section 1.2 with [LMO24, Proposition 9.19]. Suppose we are given a compact metric space $\mathcal{X}$ and a finite subset $A=\{a_{1},\ldots,a_{m}\}\subset\mathcal{X}$ with $\operatorname{diam}(A)\leq\delta$ and $\mathbf{d}^{\mathcal{X}}_{\mathrm{H}}(\mathcal{X},A)\leq\delta$. In other words, the Katz spread of $\mathcal{X}$ does not exceed $\delta$:
@@
Embed $\mathcal{X}$ in $L^{\infty}(\mathcal{X})$ via the Kuratowski map $\kappa$, and note that the set $\kappa(A)\subset L^{\infty}(\mathcal{X})$ is $\delta/2$-übercontractible. Hence, we can apply Section 1.2 and conclude that Čech lifespans of $\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X})$ do not exceed $2\delta$. Therefore, VR lifespans of $\mathcal{X}$ do not exceed $4\delta$. This is weaker than the conclusion of [LMO24, Proposition 9.19], which tells us that in this situation VR lifespans of $\mathcal{X}$ do not exceed $\operatorname{spread}(\mathcal{X})\leq\delta$.
In general, the estimate in Section 1.2 cannot be improved (see Figure 3). However, the factor of $2$ in that estimate can sometimes be removed if the comparison set $\mathcal{T}$ (i.e. the “core”) admits a $1$-Lipschitz nearest-point projection from $\mathbf{Y}$.
@@
Let $\mathcal{X}$ be a compact subset of a Banach space $\mathbf{Y}$. Assume that $\mathrm{conv}(\mathcal{X})$ admits a $1$-Lipschitz nearest-point projection to a closed convex set $\mathcal{T}\subset\mathbf{Y}$ [^15] For example, this is true whenever $\mathbf{Y}$ is Euclidean. Then any $\omega\in\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$, $k\geq 0$, has lifespan at most $\mathbf{d}_{\mathrm{H}}^{\mathbf{Y}}(\mathcal{X},\mathcal{T})$.
Remark 4.7.
The above theorem cannot be derived from the stability results in 2.15 and 2.16, as those results apply exclusively to homology classes associated with points in the persistence diagram $\widecheck{\mathrm{PD}}_{k}(\mathcal{X}\subset\mathbf{Y})$. In contrast, 4.6 applies to all elements of $\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$, which, as expressed by 2.21 and 2.22, generally includes strictly more homology classes than those represented by points in the persistence diagram $\widecheck{\mathrm{PD}}_{k}(\mathcal{X}\subset\mathbf{Y})$. Nevertheless, the proof of 4.6 shares ideas with (certain simplicial-level) proofs of the stability results.
Proof of 4.6.
Let $\delta=\mathbf{d}_{\mathrm{H}}^{\mathbf{Y}}(\mathcal{X},\mathcal{T})$. By assumption, there exists a $1$-Lipschitz nearest-point projection $\pi:\mathrm{conv}(\mathcal{X})\rightarrow\mathcal{T}$; there is also a (possibly discontinuous) map $j:\mathcal{T}\to\mathcal{X}$ such that
@@
${\mathcal{C}_{\epsilon}(\mathcal{X}\subset\mathbf{Y})}$${\mathcal{C}_{\epsilon+\delta}(\mathcal{X}\subset\mathbf{Y})}$${\mathcal{C}_{\epsilon}(\mathcal{T}\subset\mathbf{Y})}$$\scriptstyle{\pi_{\epsilon}}$$\scriptstyle{i_{\epsilon,\epsilon+\delta}}$$\scriptstyle{j_{\epsilon}}$
@@
Now, going back to the proof of the theorem: apply the homology functor $\widetilde{\mathrm{H}}_{k}$ to the diagram above and obtain the following commutative diagram:
${\widetilde{\mathrm{H}}_{k}(\mathcal{C}_{\epsilon}(\mathcal{X}\subset\mathbf{Y}))}$${\widetilde{\mathrm{H}}_{k}(\mathcal{C}_{\epsilon+\delta}(\mathcal{X}\subset\mathbf{Y}))}$${\widetilde{\mathrm{H}}_{k}(\mathcal{C}_{\epsilon}(\mathcal{T}\subset\mathbf{Y}))}$$\scriptstyle{\widetilde{\mathrm{H}}_{k}(\pi_{\epsilon})}$$\scriptstyle{\widetilde{\mathrm{H}}_{k}(i_{\epsilon,\epsilon+\delta})}$$\scriptstyle{\widetilde{\mathrm{H}}_{k}(j_{\epsilon})}$
Now, $\widetilde{\mathrm{H}}_{k}(\mathcal{C}_{\epsilon}(\mathcal{T}\subset\mathbf{Y}))=0$ since $\mathcal{T}$ is convex. Then,
$\widetilde{\mathrm{H}}_{k}(i_{\epsilon,\epsilon+\delta})=\widetilde{\mathrm{H}}_{k}(j_{\epsilon})\circ\widetilde{\mathrm{H}}_{k}(\pi_{\epsilon})=0,$